# Understanding and Mitigating Compositional Issues in Text-to-Image Generative Models

### v1

@@ Abstract
Recent text-to-image diffusion-based generative models have the stunning ability to generate highly detailed and photo-realistic images and achieve state-of-the-art low FID scores on challenging image generation benchmarks. However, one of the primary failure modes of these text-to-image generative models is in composing attributes, objects, and their associated relationships accurately into an image. In our paper, we investigate this compositionality-based failure mode and highlight that imperfect text conditioning with CLIP text-encoder is one of the primary reasons behind the inability of these models to generate high-fidelity compositional scenes. In particular, we show that (i) there exists an optimal text-embedding space that can generate highly coherent compositional scenes which shows that the output space of the CLIP text-encoder is sub-optimal, and (ii) we observe that the final token embeddings in CLIP are erroneous as they often include attention contributions from unrelated tokens in compositional prompts. Our main finding shows that the best compositional improvements can be achieved (without harming the model’s FID scores) by fine-tuning only a simple linear projection on CLIP’s representation space in Stable-Diffusion variants using a small set of compositional image-text pairs. This result demonstrates that the sub-optimality of the CLIP’s output space is a major error source. We also show that re-weighting the erroneous attention contributions in CLIP can also lead to improved compositional performances, however these improvements are often less significant than those achieved by solely learning a linear projection head, highlighting erroneous attentions to be only a minor error source. The code is available and can be accessed at https://github.com/ArmanZarei/Mitigating-T2I-Comp-Issues.
Figure 1: Overview of our analysis and proposed methods. The figure identifies two sources of errors in Stable Diffusion’s inability to generate compositional prompts: (i) erroneous attention contribution in CLIP (minor) and (ii) sub-optimal CLIP text embedding (major). We propose a window-based linear projection (WiCLP) to address these issues, applying a linear projection to a window around each token to improve embeddings.
@@ 1 Introduction
Text-to-image diffusion-based generative models [21, 18, 19, 22] have achieved photo-realistic image generation capabilities on user-defined text prompts. However recent works [12] have designed compositionality benchmarks to show that these text-to-image models have low fidelity to simple compositionality prompts such as those consisting of attributes, objects, and their associated relations (e.g., “a red book and a yellow vase”). This hinders the use of these generative models in various creative scenarios where the end-user wants to generate a scene where the composition is derived from words (and their relationships) in the prompt.
@@
Existing works [4, 9, 1, 26] propose various ways to improve compositionality in text-to-image models. These works primarily rely on modifying the cross-attention maps by leveraging bounding box annotations and performing a small optimization in the latent space during inference. Recent methods based on fine-tuning [12] the UNet also lead to improved compositonality. Despite the progress, the core reasons behind compositionality failures in text-to-image models remain unclear. Understanding these reasons helps designing effective methods that can augment text-to-image models with improved compositional capabilities.
@@
In our paper, we investigate possible reasons behind compositionality failures in text-to-image generative models. We identify two sources of errors: (i) We observe that output token embeddings in CLIP have significant attention contributions from irrelevant tokens, thereby introducing errors in generation. We then compare the internal attention contributions in CLIP for compositional prompts to the T5 text-encoder which has been shown to display strong compositional capabilities in DeepFloyd [^1] https://huggingface.co/DeepFloyd/IF-I-M-v1.0. We quantitatively find that the T5 text-encoder displays significantly lesser erroneous attention contributions than CLIP, highlighting a potential reason towards its improved compositionality. (ii) Sub-optimality of CLIP output space on compositional prompts: We observe that optimizing the text embeddings, while utilizing a frozen Stable-Diffusion UNet, effectively generates images with compositional scenes. We find out that there exists a text-embedding space capable of generating highly coherent images with compositional scenes for various attributes (e.g., color, texture, shape) which highlights that the existing CLIP output space is sub-optimal. These results indicate that the output space of the CLIP text-encoder could be further improved to enable text-to-image models to generate more accurate compositional scenes.
Leveraging our observations on the deficiencies of the CLIP output space, we show that we can improve the output space of the CLIP text-encoder to better align with the optimal space by applying a linear projection on top of CLIP (see Figure  1). This leads to stronger compositional performances. In particular, we propose Window-based Compositional Linear Projection (WiCLP), a lightweight fine-tuning method that significantly enhances the model’s performance on compositional prompts, achieving results comparable to other existing baselines (see Figure 2), while maintaining the model’s clean accuracy, as indicated by a low FID on clean prompts. We also show that reweighting the erroneous attention contributions in CLIP can lead to improved compositional performances, however, the improvements often lag behind WiCLP. This result shows that the sub-optimal alignment of the CLIP text-encoder to the UNet is a major error source compared to erroneous attention contributions in CLIP.
Fine-tuning a subset of components of the diffusion model can result in an increase in the FID score for clean prompts. While fine-tuning only a linear projection partially mitigates this, we find that applying it over all the time steps results in an increase in FID. To mitigate this, we introduce Switch-Off where we only apply WiCLP during the initial steps of generation, switching it off for the remaining steps. This enables the model to obtain a coherent compositional scene in early steps (crucial for compositional prompts) while retaining clean accuracy on surrounding prompts, as the generation in final steps is guided by the original text-encoder not the augmented one that maps to the optimized space.
@@
  • We perform an in-depth analysis of the reasons behind compositionality failures in open-source text-to-image generative models, highlighting two reasons for them.
  • Leveraging our observations, we propose WiCLP for Stable Diffusion v-1.4 and v-2 which can augment the models with improved compositionality while preserving their clean accuracy on surrounding prompts. We observe improvements of $16.18\%,15.15\%$, and $9.51\%$ on SD v1.4 and $14.35\%,11.14\%$, and $6\%$ on SD v2 in VQA scores [12] across color, texture, and shape datasets, respectively. Our method achieves competitive VQA scores compared to other baselines while having better FID on clean prompts.
Overall, our paper provides quantitative evidence elucidating the compositional challenges within text-to-image models and strong baselines to mitigate such issues.
SD v1.4 CLP SD v2 WiCLP
A blue backpack and a red chair Refer to caption Refer to caption Refer to caption Refer to caption
A yellow book and a red vase Refer to caption Refer to caption Refer to caption Refer to caption
Figure 2: Qualitative comparison between CLP and WiCLP vs the baselines
@@ 2 Background
A recent work Huang et al. [12] introduces a benchmark for testing compositionality in text-to-image models showing the susceptibility of open-source text-to-image models on simple compositional prompts. In addition, the authors also propose a fine-tuning baseline to augment text-to-image models with improved compositionality. The compositionality issue can also be addressed at inference time by modifying the cross-attention maps leveraging hand-crafted loss functions and bounding boxes generated from a language model [4, 9, 1, 26, 17, 13, 15]. However, Huang et al. [12] show that a data-driven and fine-tuning approach is more suitable towards improving compositionality in text-to-image models. Our paper specifically targets understanding the source of compositionality errors in text-to-image models which is one of the open research questions in this area.
Interpretability of Text-to-Image Generative Models.
There have been recent efforts to interpret text-to-image models like Stable Diffusion. DAAM [23, 10] studies the generation process in diffusion models by analyzing cross-attention maps between text tokens and image pixels, highlighting their semantic precision. Basu et al. [2] use causal tracing to understand how knowledge is stored in models like Stable Diffusion v1 while [3] propose a mechanistic approach to localize knowledge in cross-attention layers of various text-to-image models. Chefer et al. [5] explore concept decomposition in diffusion models.
2.1 Text-to-image Diffusion Models: Training and Inference
In diffusion models, noise is added to the data following a Markov chain across multiple time-steps $t\in[0,T]$. Starting from an initial random real image $\mathbf{x}_{0}$ along with its caption $c$, $(\mathbf{x}_{0},c)\sim\mathcal{D}$, the noisy image at time-step $t$ is defined as $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{(1-\alpha_{t})}\mathbf{\epsilon}$. The denoising network denoted by $\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c},t)$ is pre-trained to denoise the noisy image $\mathbf{x}_{t}$ to obtain $\mathbf{x}_{t-1}$. For better training efficiency, the noising along with the denoising operation occurs in a latent space defined by $\mathbf{z}=\mathcal{E}(\mathbf{x})$, where $\mathcal{E}$ is an encoder such as VQ-VAE [24]. Usually, the conditional input $\mathbf{c}$ to the denoising network $\epsilon_{\theta}(.)$ is a text-embedding of the caption $c$ through a text-encoder $\mathbf{c}=v_{\gamma}(c)$. The pre-training objective for diffusion models can be defined as follows:
@@ 2.1 Text-to-image Diffusion Models: Training and Inference
where $\theta$ is the set of learnable parameters in the UNet $\epsilon_{\theta}$. During inference, where the objective is to synthesize an image given a text-embedding $\mathbf{c}$, a random Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,I)$ is iteratively denoised for a fixed range of time-steps to produce the final image.
2.2 Dataset Collection
We utilize the T2I-CompBench dataset [12], focusing on three categories: color, texture, and shape, each with 1000 prompts. To generate high-quality images, we use three generative models: SD 1.4 [21], DeepFloyd, and SynGen [20], creating 100 samples per prompt with SD 1.4, 60 with DeepFloyd, and 50 with SynGen. This ensures a wide variety of generated images, leveraging each model’s strengths. We focus on the disentangled BLIP-Visual Question Answering (VQA) score proposed by Huang et al. [12] as a metric for evaluating image quality. The VQA score assesses how accurately an image represents the compositional elements described in a prompt and is more closely correlated with human assessment than other metrics like CLIP-Score [11]. For each prompt, we combined all 210 samples from the three models and selected the top 30 with the highest VQA scores, ensuring the final dataset consisted of images that most accurately reflected the prompts.
3 Source (i) : Erroneous Attention Contributions in CLIP
[…]
In this Section, we leverage attention contributions [7, 6] to analyze the text-embeddings of compositional prompts in the CLIP text-encoder (which is commonly used in many text-to-image models) and compare them with T5-text encoder of DeepFloyd, a model which results in stronger compositionality. Many of the compositional prompts from Huang et al. [12] have a decomposable template of the form $\mathbf{a}_{i}\ \mathbf{o}_{j}+\mathbf{a}_{j}\ \mathbf{o}_{j}$, where $\mathbf{a}_{i},\mathbf{a}_{j}$ are attributes (e.g., “black”, “matted”) while $\mathbf{o}_{i},\mathbf{o}_{j}$ describe objects (e.g., “car”, “bag”). We use attention contributions to understand how the text-embeddings of the compositional tokens (e.g., $\mathbf{a}_{i},\mathbf{a}_{j},\mathbf{o}_{i},\mathbf{o}_{j}$) are formed for both T5 and CLIP over the layers of these models.
The attention mechanism in layer $\ell$ of a transformer consists of four weight matrices $W_{\text{q}},W_{\text{v}},W_{\text{k}},W_{\text{o}}$ [25]. Each of these weight matrices is divided into $H$ heads denoted by $W_{\text{q}}^{h},W_{\text{v}}^{h},W_{\text{k}}^{h}\in\mathbb{R}^{d\times d_{h}},W_{\text{o}}^{h}\in\mathbb{R}^{d_{h}\times d}$ for all $h\in\left[H\right]$. Note that $d_{h}$ is the dimension of the internal token embeddings. We omit $\ell$ for simplicity, but each layer has its own attention matrices. These matrices are applied on the token embeddings of the output of layer $\ell-1$, denoted by $\bar{\mathbf{x}}_{j}$ for token $j$ in that layer. We denote by $\text{q}_{j}^{h},\text{k}_{j}^{h}$, and $\text{v}_{j}^{h}$ the projection of $\bar{\mathbf{x}}_{j}$ on query, key, and value matrices of the $h$-th head of layer $\ell$. More precisely,
@@ 3 Source (i) : Erroneous Attention Contributions in CLIP
Notably, $\text{cont}_{i,j}$ is a significant metric that quantifies the contribution of a token $j$ to the norm of a token $i$ at layer $\ell$. We employ this metric to identify layers in which important tokens highly attend to unintended tokens, or lowly attend to intended ones. We refer to Appendix C.1 for more details on attention contribution.
3.1 Key Finding: T5 has less erroneous attention contributions than CLIP
@@
We refer to Figure 3 that visualizes attention contribution of both T5 and CLIP text-encoder in the last layer $(\ell=11)$ for the prompt "a green bench and a red car". Ideally, the attention mechanism should guide the token "car" to focus more on "red" than "green", but in the last layer of the CLIP text-encoder, "car" significantly attends to "green". In contrast, T5 shows a more consistent attention pattern, with "red" contributing more to the token "car" and "green" contributing more to the token "bench".
@@
We further conduct an extensive analysis on specific types of prompts, consisting of $780$ prompts of color dataset and $582$ prompts of texture dataset, each structured as “$\mathbf{a}_{1}$ $\mathbf{o}_{1}$ and $\mathbf{a}_{2}$ $\mathbf{o}_{2}$.” For each prompt, we obtain attention contributions in all layers and count the number of layers where unintended attention contributions occur. In the CLIP text-encoder, unintended attention occurs when $\mathbf{o}_{2}$ attends more to $\mathbf{a}_{1}$ than $\mathbf{a}_{2}$. For T5, it occurs when $\mathbf{o}_{2}$ attends more to $\mathbf{a}_{1}$ than $\mathbf{a}_{2}$, or $\mathbf{o}_{1}$ attends more to $\mathbf{a}_{2}$ than $\mathbf{a}_{1}$. Figure 3 quantitatively compares unintended attention on various prompts between CLIP text-encoder and T5. The T5 model shows improvement in our metric over the CLIP text-encoder, supporting the hypothesis that erroneous attention mechanisms in CLIP may contribute to the poor compositionality performance of CLIP-based text-to-image models.
Figure 3: Comparative analysis of VQA Scores between CLIP text-embeddings and optimized text-embeddings using Stable Diffusion v1.4 across color, texture, and shape categories. Results show CLIP text embeddings achieved scores of 0.3615 for color, 0.4306 for texture, and 0.3619 for shape, while optimized text embeddings achieved scores of 0.7513 for color, 0.7254 for texture, and 0.58728 for shape.
3.2 Zero-shot Attention Reweighting
Inspired by attention mechanism shortcomings of CLIP text-encoder, we aim to improve compositionality of CLIP-based diffusion models by zero-shot reweighting of the attention maps. Specifically, we apply a hand-crafted zero-shot manipulation of the attention maps in certain layers of the CLIP text-encoder to effectively reduce unintended attentions while enhancing meaningful ones. This zero-shot reweighting is applied to the logits before the Softmax layer in the last three layers of the text-encoder. To be more precise, we compute a matrix $M\in\mathbb{R}^{n\times n}$ and add this matrix to the logits of attention mechanism. For each head $h$, the new attention values are obtained as follows and then propagated to the subsequent layers of the text-encoder:
We set the values in $M$ by considering the ideal case where no incorrect attentions occur in the mechanism. For example, for prompt “a green bench and a red car", we ensure that the token "car" does not attend to the token "green" by assigning a sufficiently large negative value to the corresponding entry in matrix $M$. Further details on how we obtain matrix $M$ can be found in Appendix C.2.
Key Results. Applying zero-shot attention reweighting with matrix $M$ on 780 compositional prompts of color dataset, we achieved a $2.93\%$ improvement in VQA Scores. We provide examples of effective zero-shot reweighting, demonstrating its impact on mitigating compositionality issues in Appendix C.2. From this result, we can infer that although erroneous attention contributions in the CLIP text-encoder is one source of error, it is not the primary error source due to modest improvements in compositional accuracy. In the next section, we investigate the sub-optimality of the output space of CLIP text-encoder, which we find to be a significant source of error.
4 Source (ii) : Sub-optimality of CLIP Text-Encoder for Compositional Prompts

### v2

# Improving Compositional Attribute Binding in Text-to-Image Generative Models via Enhanced Text Embeddings

@@ Abstract
Text-to-image diffusion-based generative models have the stunning ability to generate photo-realistic images and achieve state-of-the-art low FID scores on challenging image generation benchmarks. However, one of the primary failure modes of these text-to-image generative models is in composing attributes, objects, and their associated relationships accurately into an image. In our paper, we investigate compositional attribute binding failures, where the model fails to correctly associate descriptive attributes (such as color, shape, or texture) with the corresponding objects in the generated images, and highlight that imperfect text conditioning with CLIP text-encoder is one of the primary reasons behind the inability of these models to generate high-fidelity compositional scenes. In particular, we show that (i) there exists an optimal text-embedding space that can generate highly coherent compositional scenes showing that the output space of the CLIP text-encoder is sub-optimal, and (ii) the final token embeddings in CLIP are erroneous as they often include attention contributions from unrelated tokens in compositional prompts. Our main finding shows that significant compositional improvements can be achieved (without harming the model’s FID score) by fine-tuning only a simple and parameter-efficient linear projection on CLIP’s representation space in Stable-Diffusion variants using a small set of compositional image-text pairs.
Improving Compositional Attribute Binding in Text-to-Image Generative Models via Enhanced Text Embeddings
Arman Zareithanks: Equal contribution. Correspondence to: azarei@umd.edu, krezaei@umd.edu, Keivan Rezaei11footnotemark: 1, Samyadeep Basu, Mehrdad Saberi, Mazda Moayeri, Priyatham Kattakinda, Soheil Feizi Department of Computer Science University of University of Maryland
Figure 1: Overview of our analysis and proposed methods. The figure identifies two sources of errors in Stable Diffusion’s inability to generate compositional prompts: (i) erroneous attention contribution in CLIP (minor) and (ii) sub-optimal CLIP text embedding (major). We propose a window-based linear projection (WiCLP), applying linear projection to a token’s surrounding window to enhance embeddings.
@@ 1 Introduction
Text-to-image diffusion-based generative models (Rombach et al., 2021; Podell et al., 2023; Ramesh et al., 2021; Saharia et al., 2022) have achieved photo-realistic image generation capabilities on user-defined text prompts. However, recent studies (Huang et al., 2023) reveal that text-to-image models struggle with maintaining high fidelity when handling simple compositional prompts, such as those consisting of attributes, objects, and their associated relations (e.g., “a red book and a yellow vase”). This hinders the use of these generative models in various creative scenarios where the end-user wants to generate scenes that accurately reflect the composition and relationships specified in the prompt.
@@
Existing approaches (Chefer et al., 2023; Feng et al., 2023; Agarwal et al., 2023; Wang et al., 2023) explore various strategies to enhance compositionality in text-to-image models. These methods primarily focus on modifying cross-attention maps by utilizing bounding box annotations and performing optimizations in the latent space during inference. Recent advancements, such as fine-tuning the UNet (Huang et al., 2023), have also demonstrated improvements in compositionality. However, the core reasons behind compositionality failures remain poorly understood. Gaining insights into these root causes is crucial for developing more effective approaches to augment these models with enhanced compositional capabilities.
@@
In our paper, we investigate the potential causes of compositional attribute binding failures in text-to-image generative models, where the model fails to correctly associate descriptive attributes (such as color, shape, or texture) with the corresponding objects in the generated images. We identify two key sources of error: (i) Erroneous attention contributions in CLIP output token embeddings: We observe that output token embeddings in CLIP have significant attention contributions from irrelevant tokens, thereby introducing errors in generation. To explore this, we compare the internal attention contributions in CLIP for compositional prompts with the T5 text encoder, known for its stronger compositionality. Quantitative analysis shows that T5 exhibits fewer erroneous attention contributions than CLIP, indicating a potential reason for its superior compositionality. (ii) Sub-optimality of CLIP output space for compositional prompts: We find out that there exists an alternative text-embedding space capable of generating highly coherent images from compositional prompts. This indicates that the current CLIP output space is inherently sub-optimal. Specifically, optimizing CLIP’s text embeddings, while keeping the Stable Diffusion UNet frozen, converges to a more effective embedding space, enabling better compositional image generation. These findings highlight that refining the output space of the CLIP text encoder could play a critical role in enhancing compositionality.
Building on our observations about the deficiencies of CLIP and identifying its text-embedding space as a core issue in compositional attribute binding, we explore augmenting diffusion models with a lightweight module to enhance the text-encoder’s output and improve compositionality. Remarkably, a simple linear projection achieves significant improvements, comparable or superior to full fine-tuning of CLIP or training more complex networks on top of it. We demonstrate that this linear projection effectively aligns the CLIP text-encoder’s output with a more optimal embedding space (see Figure 1), leading to significantly stronger compositional performances.
In particular, we introduce Window-based Compositional Linear Projection (WiCLP), a lightweight fine-tuning method that significantly improves the model’s performance on compositional prompts (Figure 2), achieving results comparable to or surpassing existing methods. Additionally, WiCLP preserves the model’s overall performance, maintaining high fidelity on clean prompts as evidenced by a low FID score, while offering a solution that is both parameter efficient and speed efficient. This ensures robust compositional capabilities without compromising the model’s general effectiveness.
@@
  • We perform an in-depth analysis of the reasons behind compositionality failures in text-to-image generative models, with a particular focus on investigating the attribute binding aspect of compositionality. We highlight two key reasons contributing to these failures.
  • Building on our observations, we propose WiCLP as an enhancement for SD v1.4, SD v2, SDXL, SD v3, DeepFloyd, and PixArt-$\alpha$. This method significantly improves the models’ compositional attribute binding, while preserving accuracy on standard prompts. We observe improvements of $16.18\%,15.15\%$, and $9.51\%$ on SD v1.4, $14.35\%,11.14\%$, and $6\%$ on SD v2, $20.31\%,13.4\%$, and $5\%$ on SDXL, and $14.16\%$, $9.82\%$, and $2.63\%$ on PixArt-$\alpha$ in VQA scores across color, texture, and shape datasets, respectively. Our method outperforms or matches existing baselines in VQA scores, while achieving a superior FID score. It requires fewer parameters for optimization and enables faster inference, making it both efficient and effective.
SD v1.4 WiCLP SD v2 WiCLP
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A blue backpack and a red chair
SDXL WiCLP PixArt-$\alpha$ WiCLP
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
A yellow book and a red vase
Figure 2: Qualitative comparison of baselines and our projection method (WiCLP). Incorporating WiCLP significantly improves image alignment with the prompts.
@@ 2 Background
Compositionality in text-to-image models refers to the ability of a model to accurately capture the correct compositions of objects, their corresponding attributes, and the relationships between objects described in a given prompt. Huang et al. (2023) introduced a benchmark designed to evaluate compositionality in text-to-image models, highlighting the limitations of models when handling compositional prompts. The benchmark employs disentangled BLIP-Visual Question Answering (VQA) as a metric for assessing image compositional quality. The VQA score assesses how accurately an image captures the compositional elements described in the prompt by utilizing a vision-language model. This metrics demonstrates a closer correlation with human judgment compared to metrics like CLIP-Score (Hessel et al., 2021). The authors also proposed a fine-tuning baseline to enhance compositionality in these models. Alternatively, compositionality issues can be addressed at inference by modifying cross-attention maps using hand-crafted loss functions and bounding boxes derived from a language model (Chefer et al., 2023; Feng et al., 2023; Agarwal et al., 2023; Wang et al., 2023; Nie et al., 2024; Lian et al., 2023; Liu et al., 2022a). However, Huang et al. (2023) showed that data-driven fine-tuning is more effective for improving compositionality.
Text-to-image Diffusion Models
In diffusion models, noise is added to the data following a Markov chain across multiple time-steps $t\in[0,T]$. Starting from an initial random real image $\mathbf{x}_{0}$ along with its caption $c$, $(\mathbf{x}_{0},c)\sim\mathcal{D}$, the noisy image at time-step $t$ is defined as $\mathbf{x}_{t}=\sqrt{\alpha_{t}}\mathbf{x}_{0}+\sqrt{(1-\alpha_{t})}\mathbf{\epsilon}$. The denoising network denoted by $\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c},t)$ is trained to denoise the noisy image $\mathbf{x}_{t}$ to obtain $\mathbf{x}_{t-1}$. For efficiency, the noising and the denoising operations occur in a latent space defined by $\mathbf{z}=\mathcal{E}(\mathbf{x})$, where $\mathcal{E}$ is an encoder such as VQ-VAE (van den Oord et al., 2017). Usually, the conditional input $\mathbf{c}$ to the denoising network $\epsilon_{\theta}(.)$ is a text-embedding of the caption $c$ through a text-encoder $\mathbf{c}=v_{\gamma}(c)$. The training objective for diffusion models can be defined as follows:
@@
where $\theta$ is the set of learnable parameters in the UNet $\epsilon_{\theta}$. During inference, given a text-embedding $\mathbf{c}$, a random Gaussian noise $\mathbf{z}_{T}\sim\mathcal{N}(0,I)$ is iteratively denoised to produce the final image.
3 Sources of Compositionality Failures
This section conducts an in-depth analysis of compositional attribute binding failures in text-to-image models, focusing on the CLIP text-encoder.
3.1 Source (i) : Erroneous Attention Contributions in CLIP
Figure 3: The heatmap illustrates unintended attention contributions in CLIP, while highlighting the more accurate performance of T5.
Figure 4: Quantitatively, we find CLIP to have significantly higher erroneous attention contributions averaged across prompts of color and texture datasets.
In this section, we leverage attention contributions (Elhage et al., 2021; Dar et al., 2023) to analyze how the final text-embeddings of compositional prompts are obtained by the CLIP text-encoder, a widely adopted component in many text-to-image models. We then compare these attention contribution patterns with those produced by the T5 text-encoder which is known for its stronger compositional capabilities. Many of the compositional prompts from Huang et al. (2023) have a decomposable template of the form $\mathbf{a}_{i}\ \mathbf{o}_{j}+\mathbf{a}_{j}\ \mathbf{o}_{j}$, where $\mathbf{a}_{i},\mathbf{a}_{j}$ are attributes (e.g., “black”, “matted”) and $\mathbf{o}_{i},\mathbf{o}_{j}$ represent objects (e.g., “car”, “bag”).
The attention mechanism in layer $\ell$ of a transformer consists of four weight matrices $W_{\text{q}},W_{\text{v}},W_{\text{k}},$ and $W_{\text{o}}$ (Vaswani et al., 2017). Each of these matrices is divided into $H$ heads, denoted by $W_{\text{q}}^{h},W_{\text{v}}^{h},W_{\text{k}}^{h}\in\mathbb{R}^{d\times d_{h}},W_{\text{o}}^{h}\in\mathbb{R}^{d_{h}\times d}$, where $h\in\left[H\right]$. Here, $d_{h}$ denotes the dimensionality of the internal token embeddings. For simplicity, we omit $\ell$, but each layer has its own attention matrices. These matrices operate on the token embeddings produced by the previous layer ($\ell-1$), denoted as $\bar{\mathbf{x}}_{j}$ for token $j$. We further denote the projections of $\bar{\mathbf{x}}_{j}$ onto the query, key, and value matrices of the $h$-th attention head in layer $\ell$ as $\text{q}_{j}^{h}$, $\text{k}_{j}^{h}$, and $\text{v}_{j}^{h}$, respectively. More precisely,
@@ 3.1 Source (i) : Erroneous Attention Contributions in CLIP
Notably, $\text{cont}_{i,j}$ is a metric that quantifies the contribution of a token $j$ to the norm of a token $i$ at layer $\ell$. We employ this metric to identify layers in which important tokens highly attend to unintended tokens, or lowly attend to intended ones. See Appendix B.1 for details on attention contribution.
@@
Key Finding: T5 has less erroneous attention contributions than CLIP. We refer to Figure 4 that visualizes attention contribution of both T5 and CLIP text-encoder in the last layer $(\ell=11)$ for the prompt “a green bench and a red car". Ideally, the attention mechanism should guide the token “car” to focus more on “red” than “green”, but in the last layer of the CLIP text-encoder, “car” significantly attends to “green”. In contrast, T5 shows a more consistent attention pattern, with “red” contributing more to the token “car” and “green” contributing more to the token “bench”.
@@
We further conduct a comprehensive analysis focusing on specific types of compositional prompts from the T2I-CompBench dataset (Huang et al., 2023). This includes $780$ prompts from the color category and $582$ prompts from the texture category of this dataset, each following the structured format: “$\mathbf{a}_{1}$ $\mathbf{o}_{1}$ and $\mathbf{a}_{2}$ $\mathbf{o}_{2}$”. For each prompt, we obtain attention contributions in all layers and count the number of layers where unintended attention contributions occur. In the CLIP text-encoder, unintended attention occurs when $\mathbf{o}_{2}$ attends more to $\mathbf{a}_{1}$ than $\mathbf{a}_{2}$. For T5, it occurs when $\mathbf{o}_{2}$ attends more to $\mathbf{a}_{1}$ than $\mathbf{a}_{2}$, or $\mathbf{o}_{1}$ attends more to $\mathbf{a}_{2}$ than $\mathbf{a}_{1}$. Figure 4 provides a quantitative comparison of unintended attention across various prompts between the CLIP text-encoder and T5. The T5 model demonstrates superior performance on our metric compared to the CLIP text-encoder, reinforcing the hypothesis that erroneous attention mechanisms in CLIP may contribute to its weaker compositional performance in text-to-image models. Additional details can be found in Appendix B.4. Further experiments with other text-encoders are also reported in Appendix B.3.
To address the attention shortcomings of the CLIP text-encoder, we explored zero-shot reweighting of attention maps in CLIP to reduce unintended attentions while enhancing meaningful ones. While this improved baseline performance, it fell short of our primary method discussed in the following sections. See Appendix B.2 for more details.
3.2 Source (ii) : Sub-optimality of CLIP Text-Encoder for Compositional Prompts
Figure 5: Sub-optimality of CLIP Text-Encoder for Compositional Prompts. Optimizing a learnable vector to represent an improved text embedding, while keeping the UNet frozen, enables the generation of more compositionally accurate images.